OpenStudy (anonymous):

The average amount of water flowing over Niagara Falls is 1.68 X 10^5 m^3/min. a. How many cubic meters of water flow over the falls in one day? b. How many cubic meters of water flow over the falls in one year?

jimthompson5910 (jim_thompson5910):

1 hour = 60 min 24*1 hour = 24*60 min 24 hours = 1440 min 1 day = 1440 min

jimthompson5910 (jim_thompson5910):

a) (1.68 X 10^5 m^3/min) * (1440 min/1 day) = 1.68 X 10^5*1440 = 241,920,000 m^3/day which converts to 2.42 x 10^8 in scientific notation So roughly 2.42 x 10^8 cubic meters of water flow over the falls in one day